OpenStudy (anonymous):

solution to the equation 8^(y + 3) = 2^y ? is y= -4.5 correct?

OpenStudy (anonymous):

write as \[\Large 2^{3(y+3)}=2^y\]

OpenStudy (anonymous):

answer is either +/- 4.5 or +/-1.5 but 4.5 is wrong and i know it will be on the test

OpenStudy (anonymous):

since bases are same just put then equal \[\Large 3(y+3)=y\] solve for y

OpenStudy (anonymous):

3y+9=y

OpenStudy (anonymous):

-9/4

OpenStudy (anonymous):

-4.5

OpenStudy (anonymous):

ty